OpenStudy (anonymous):

Sylvia has a total of 16 trading cards: 8 superhero cards, 4 princess cards, 2 villain cards, and 2 wildlife cards. What is the ratio of princess cards to the total number of trading cards in simplest form? to

OpenStudy (anonymous):

____to____

OpenStudy (anonymous):

@Greg_D

OpenStudy (anonymous):

@Surana

OpenStudy (anonymous):

the ratio is tue quotient of the number of princess card divided by the total number of cards...

OpenStudy (anonymous):

i think 2 to 2

OpenStudy (anonymous):

how many princess cards are there? and how many cards in total?

OpenStudy (anonymous):

4/2=2 and there r 2wildlife

OpenStudy (anonymous):

4 p cards 32 cards in all 32/4=8

OpenStudy (anonymous):

8+4+2+2= d 16 = d 4/16= 1/4 sorry read that wrong at first

OpenStudy (anonymous):

"Sylvia has a total of 16 trading cards" of which : superhero: 8 princess: 4 villain: 2 wildlife: 2 so, the ratio of princess to total is 4 to 16 but that is not the simplest form... can you simplify 4/16 ?

OpenStudy (anonymous):

so its 4 to 16

OpenStudy (anonymous):

yes, but you are asked to give the simplest form that means the minimal equivalent fraction what if you divide both numerator and denominator by 4 ?

OpenStudy (anonymous):

4/16 simplify's into 1/4

OpenStudy (anonymous):

by 4

OpenStudy (anonymous):

so is it 4 to 2

OpenStudy (anonymous):

4 to 16 or 1 to 4 .... they are the same ratio, but you are asked for the simplest... so 1 to 4 !